Red Raider Company uses a plantwide overhead rate with direct labor hours as the allocation base. Next year, 560,000 units are expected to be produced requiring 0.90 direct-labor hours each. How much overhead will be assigned to each unit produced given the following estimated amounts?

Estimated: Department 1 Department 2
Manufacturing overhead costs $ 2,562,000 $ 948,000
Direct labor hours 184,000 DLH 126,000 DLH
Machine hours 31,600 MH 9,600 MH
Answer #1
Estimated Manufacturing overhead costs 3510000 =2562000+948000
Divide by Estimated direct labor hours 310000 =184000+126000
Plantwide overhead rate 11.32
Overhead assigned to each unit 10.19 per unit =11.32*0.90